UT 20170923 LUCI1/2 AO Nighttime Commissioning

Observer LUCI-Team: JHeidt, DThompson (LBTO)
Observer AO-Team: DMiller, GTaylor (LBTO)
Telescope Operator: Sallenson

Observations done remotely from Tucson

Plan for the night

There have been several issues with the AO-systems last night, which need to be addressed, ie

- we start with AO-testing

- LUCI AO-ima (short) when the above is concluded and problems understood/solved

- LUCI AFC-ima if there is time left

Summary

No data taken, closed due to high humidity all night

Details

All times are UT
Files are luci1/2.20170923.NNNN.fits

01:20 Sunset, some high-level scattered clouds, but many low-level, fast moving ones,
T = +6.5d, hum = 90%, wind = 15-20m/s

We stay closed for the moment being

06:30 Hum 100%, wind 15-20 m/s, still closed

...closed for the rest of the night

-- %USERSIG{JochenHeidt - 2017-09-23}%
Topic revision: r4 - 23 Sep 2017, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback